OpenStudy (anonymous):

Find the slope of the tangent line to the curve √4x+1y + √2xy = 6.06 @ the point (3,1) I got this (-1-sqrt(2) sqrt(x) y)/(sqrt(x) (1+sqrt(2) x)), then I plugged in the values 3 and 1 the answer is -0.379877994777674 But for some reason I still got it wrong I guess not sure Any help?

OpenStudy (amistre64):

when more than one digit appears in the sqrt, its best to use the () to indicate its extent.

OpenStudy (amistre64):

im assuming its √(4x+1y) + √(2xy) = 6.06

OpenStudy (anonymous):

Yes....that's right

OpenStudy (anonymous):

Sorry

OpenStudy (amistre64):

√(4x+1y) + √(2xy) = 6.06 (4x'+y') 2x'y+2xy' --------- + --------- = 0 2√(4x+1y) 2√(2xy)

OpenStudy (amistre64):

4 y' 2y 2xy' --------- + --------- +---------+---------= 0 2√(4x+1y) 2√(4x+1y) 2√(2xy) 2√(2xy) now we can gather the y's and solve

OpenStudy (anonymous):

so just plug in the values or we still have to do the y right?